OpenStudy (anonymous):

inside one leaf of the four-leaved rose r = 7 sin 2(theta)

zepdrix (zepdrix):

what do you mean by `inside`? Find the area inside?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

The formula for Area in Polar is given by:\[\Large A\quad=\quad \frac{1}{2}\int\limits_a^b r^2\;d \theta\]I guess we have to be careful with our bounds though right? Cause we want to make sure we go from one end of the petal to the other. So we need to determine whether it looks like this: Or like this:Oh boy my artwork is terrible :) lol